The average

daily demand

for laptops at Best Buy is 20, with a std. deviation of 5 units.

The lead time is constant at 2 days. Find the reorder point if management wants a 90% service

level (i.e., risk stockouts only 10% of the time). How much is the safety stock?
Answer #1

Safety Stock = Z * standard deviation * sqrt(lead time)

Z = 1.28 (for 10% stockout risk)

Safety stock = 1.28 * 5 * sqrt(2) = 9.051

ROP = Daily demand * Lead time + Safety stock

= 20 * 2 + 9.051

= 49.051

Do give a positive rating if you think this answer helped